theyadoreshayyy:

Complete the square to rewrite the following equation. Identify the center and radius of the circle. You must show all work and calculations to receive credit. x2 + 2x + y2 + 4y = 20

theyadoreshayyy:

@snowflake0531

theyadoreshayyy:

@darkknight

snowflake0531:

One question, do you know how to do this at all? have you done similar problems?

theyadoreshayyy:

When i did my DBA, but she was rushing and she didnt listen when i told her i didnt understand it.

snowflake0531:

\[x^2+2x+y^2+4x=20\\(x^2+2x) + (y^2+4x)=20\] \[(\frac{b}{2})^2\] where b is ax^2 + bx + c, split it to x^2 +2x and y^2 + 4x, the b's are 2 and 4 so it's 1 and 4 \[(x^2+2x+1) +(y^2+4x+4)=20+1+4\\(x+1)^2 + (y+2)^2=25\] \[(x-h)^2+(y-k)^2=r^2\]